OpenStudy (anonymous):

least common multiple: x²-x, x-1, x²-1

OpenStudy (nikvist):

\[x^2-x=(x-1)x\quad,\quad x-1\quad,\quad x^2-1=(x-1)(x+1)\]\[LCM(x^2-x,x-1,x^2-1)=(x-1)x(x+1)\]